OpenStudy (anonymous):

Find the area above x-axis for given function on given interval by using the Fundamental theorem of Calculus. F(x)= 1+2sinx on [0,pi/2]

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

which ftc should I use first or the second?

OpenStudy (danjs):

is this just to integrate with respect to x over that interval?

OpenStudy (danjs):

F(upper)-F(lower)

OpenStudy (anonymous):

I don't think so. I believe it has already integrated considering it uses an upper case F(x).. \[F(x)=\int\limits f(x)dx\]Probably just plug in the limits

OpenStudy (danjs):

from 0 to half pi, sin(x) is always a positive number, the function is always positive

OpenStudy (danjs):

do the integral of the function from 0 to half pi

OpenStudy (anonymous):

find the area above x-axis- (for given function on given interval)

OpenStudy (danjs):

F(x) is the given function, the area under that is the definite integral value for the interval

OpenStudy (anonymous):

I wasn't sure if the notation was specific. I was taught that it was, but I guess it may not be.

OpenStudy (danjs):

yeah that is what integrating will do, it can give you an exact area value for spaces that are curved in any way on the boundries

OpenStudy (anonymous):

it has closed brackets towards it so I guess just teach me the way u is right I will check about that with my prof tmr

OpenStudy (danjs):

you see the pictures of 5 boxes, 10 boxes, infinite boxes, to estimate the area under curves

OpenStudy (anonymous):

Yeah, I'm aware. But I was afraid that the question already presumes that it was integrated and just wants the student to plug in the limits, hence the "using fundamental theorem of calculus" part.

OpenStudy (anonymous):

If notation is specific, then \[\large \int\limits f(x)dx=F(B)-F(A)\] I feel like we already have the left hand, now we just need to plug in the limits. But I could be wrong.

OpenStudy (anonymous):

you see the pictures of 5 boxes, 10 boxes, infinite boxes, to estimate the area under curves- yes, I know that

OpenStudy (anonymous):

but what are u trying to say I am confused now

OpenStudy (danjs):

the given function is F(x) = 1 + 2sin(x), they want the area between x-axis and that function. That is what the fundamental theorem will do, you need to integrate, Just make the next F even larger. haha

OpenStudy (anonymous):

X)

OpenStudy (anonymous):

okay by using the first fundmental theorem?

OpenStudy (danjs):

\[\int\limits_{0}^{\pi/2}F(x)dx \]

OpenStudy (anonymous):

that's the second i think

OpenStudy (danjs):

oh, you talking about that derivative of an integral , other thing?

OpenStudy (anonymous):

no the ftc- I think the π/2F(x)dx u gave above is the second fundamental theo of calc

OpenStudy (anonymous):

The second fundamental theorem of calculus is: \[F(x)=\int\limits_{0}^{x}f(t)dt \implies \frac{ d }{ dx }[F(x)]=f(x)\]

OpenStudy (anonymous):

@DanJS okay what is next plz explain me in terms of math I get confused when u talk in sentences.

OpenStudy (anonymous):

X)

OpenStudy (danjs):

Integrate the function and you get x - 2*cos(x) then you take the upper bound and the lower bound and evaluate those in that x-2cos(x) take the value of the upper - value of the lower, that is the answer

OpenStudy (danjs):